How far apart the two electrons be, if the force between them equals the weight of an electron? What in case of protons?

Grade:12

2 Answers

Arun
5 years ago
Dear Abhijeet
 
k e^2 / r^2 = m g 
r = e [k / (m g)]^1/2 
then 
1) r = 1.6x10^-19 [9x10^9 / (9.1x10^-31*9.8)]^1/2 = 5.08 m [electrons] 
2) r = 1.6x10^-19 [9x10^9 / (1.67x10^-27*9.8)]^1/2 = 0.12 m [protons]
 
Regards
Arun (askIITians forum expert)
ankit singh
3 years ago
by force balance
ke2/r2=mg
thenput the values of charge fo protons and
r = 1.6x10^-19 [9x10^9 / (9.1x10^-31*9.8)]^1/2 = 5.08 m [electrons] 
2) r = 1.6x10^-19 [9x10^9 / (1.67x10^-27*9.8)]^1/2 = 0.12 m [protons]